Obtain the directional derivative for a scalar field ϕ(x, y, z)=(3x^2y-y^3z^2) at the point (1, -2, -1) in the direction Î+ ĵ +k̂
1
Expert's answer
2020-04-23T12:44:17-0400

scalar field "\phi (x,y,z)=3x^2y-y^3z^2"

point "M(1;-2;-1)"

direction "\vec{i}+\vec{j}+\vec{k}" with normalized vector "\vec{n}({\frac 1 {\sqrt3}};{\frac 1 {\sqrt3}};{\frac 1 {\sqrt3}})"

To find directional derivative we need to get scalar product "grad(\phi )\cdot \vec{n}"

"grad(\phi)=({\frac {\partial\phi} {\partial x}};{\frac {\partial\phi} {\partial y}};{\frac {\partial\phi} {\partial z}})" at point "M"

"{\frac {\partial\phi} {\partial x}}=6xy=-12"

"{\frac {\partial\phi} {\partial y}}=3x^2-3y^2z^2=-9"

"{\frac {\partial\phi} {\partial z}}=-2y^3z=-16"

Then "grad(\phi)=(-12;-9;16)"

Directional derivative equal "grad(\phi )\cdot \vec{n}=-12\cdot{\frac 1 {\sqrt3}}-9\cdot{\frac 1 {\sqrt3}}-16\cdot{\frac 1 {\sqrt3}}=-{\frac {37} {\sqrt3}}"